A solid sphere and a solid cylinder are made of the same material. If they have the same mass and radius, which one has the larger rotational inertia about its center?
(a) the sphere
(b) the cylinder
(c) They are the same

...help.
Answer #1

.for solid cylinder   I = MR2 / 2        ,FOR solid sphere I = 2 /5 MR 2​ ​ .FROMthese two we can say solid sphere has larger rotational inertia about its center   (b)